### v2

# Finiteness of the set of associated primes for local cohomology modules of ideals via properties of almost factorial rings

Abstract
We investigate the finiteness of the set of associated primes for local cohomology modules $H_{I}^{i}(J)$ of an ideal $J$ generated by an $R$-sequence, through the comparison of $H_{I}^{d+1}(J)$ and $H_{I}^{d}(R/J)$, where $d=\operatorname{depth}_{I}(R)$. The properties of almost factorial rings play a key role in enabling this comparison. Under suitable conditions, we prove that the finiteness of $\operatorname{Ass}H_{I}^{d+1}(J)$ is equivalent to that of $\operatorname{Ass}H_{I}^{d}(R/J)$. Moreover, we give a few conditions under which the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ holds for all $i$.
Keywords: Local cohomology, Associated primes
1 Introduction
Let $I$ be an ideal of a Noetherian ring $R$. Although a local cohomology module $H_{I}^{i}(R)$ is not necessarily finite over $R$, we may still ask whether the set $\operatorname{Ass}H_{I}^{i}(R)$ of associated primes of $H_{I}^{i}(R)$ is finite, as this can provide insights into the finiteness properties of local cohomology modules. Indeed, the set $\operatorname{Ass}M$ is finite for any finite $R$-module $M$. There are examples in which $\operatorname{Ass}H_{I}^{i}(R)$ is infinite : one described by Katzman[Kat] in the context of $R$ being a local algebra over a field, and another given by Singh and Swanson[SiSw, Theorem 5.4] in the context of $R$ being a UFD algebra over a field. In contrast, there are examples in which $\operatorname{Ass}H_{I}^{i}(R)$ is finite. For instance, Huneke and Sharp[HuSh, Corollary 2.3] proved that $\operatorname{Ass}H_{I}^{i}(R)$ is finite for any ideal $I$ and any $i\geq 0$ under the condition that $R$ is a regular ring of prime characteristic $p>0$, and Lyubeznik[Lyu1, Corollary 3.6(c)] proved the same result under the condition that $R$ is a regular semi-local ring containing a field.
Definition 1.1.
Let $R$ be a Noetherian ring. If $\operatorname{Ass}H_{I}^{i}(R)$ is finite for any ideal $I$ of $R$ and any $i\geq 0$, $R$ is called LC-finite.
As mentioned above, regular rings of characteristic $p>0$, and regular semi-local rings containing a field are LC-finite. See [Mar], for more examples of an LC-finite ring.
Since the ring $R$ is generated by a single element as an $R$-module, one might expect that the local cohomology module $H_{I}^{i}(M)$ of a finite $R$-module $M$ has a more complicated structure than $H_{I}^{i}(R)$. The finiteness of $\operatorname{Ass}H_{I}^{i}(M)$ holds for $i=0$ and $i=1$. For $i=0$, this follows from the fact that $H_{I}^{0}(M)$ is finite over $R$. For $i=1$, the result was shown by Brodmann and Faghani[BrFa]. For $i=2$, in the case where $M$ is an ideal, if $R$ is a locally almost factorial Noetherian normal ring, then $\operatorname{Ass}H_{I}^{2}(J)$ is finite for any ideal $I$, $J$ of $R$ [Lew, Theorem 2.1]. Under the same condition, there is an example in which $\operatorname{Ass}H_{I}^{3}(J)$ is an infinite set [Lew, Section 3]. In this example, the infiniteness of $\operatorname{Ass}H_{I}^{3}(J)$ is deduced from the infiniteness of $\operatorname{Ass}H_{I}^{2}(R/J)$ which is described in [Kat]. To be precise, in this example, $H_{I}^{2}(R/J)$ is isomorphic to a submodule of $H_{I}^{3}(J)$ and all of the associated primes of $H_{I}^{2}(R/J)$ are inherited by $H_{I}^{3}(J)$.
We investigate the finiteness of the set of associated primes for local cohomology modules of ideals, based on the finiteness of that for local cohomology modules of Noetherian rings. The background involves the following question that was studied in [Lew]:
Question 1.2.
Let $R$ be an LC-finite regular ring, $J$ an ideal of $R$ generated by an $R$-sequence, and $I$ an ideal of $R$ containing $J$. Fix $i\geq 1$. Then, $\operatorname{Ass}H_{I}^{i}(J)$ is finite if $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite?
For $i\leq 2$, it is true. If $R/J$ is a locally almost factorial normal ring, it is true for $i\leq 4$ [Lew, Theorem 4.2]. Question 1.2 is reduced to the following question, see [Lew, Lemma 4.1 and Corollary 4.3]:
Question 1.3.
Let $R$ be an LC-finite regular ring, $J$ an ideal of $R$ generated by an $R$-sequence of length at least $2$, and $I$ an ideal of $R$ containing $J$. Fix $i\geq 1$, and suppose $\operatorname{depth}_{I}(R)=i-1$. Then, $\operatorname{Ass}H_{I}^{i}(J)$ is finite if $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite?
We consider cases according to the value of $\operatorname{depth}_{I}(R/J)$, and in each case, we give an affirmative answer to the question under suitable assumptions : For the case $\operatorname{depth}_{I}(R/J)\leq 1$, the question is true if $R/J$ is a locally almost factorial normal ring (see [Lew, Theorem 4.2] or Section 3). For the case $\operatorname{depth}_{I}(R/J)\geq 2$, we give the following statement, which generalizes the case $\operatorname{depth}_{I}(R/J)=1$. This result will be proved in Section 4.
Theorem 1.4.
Let $R$ be a Noetherian ring, $J$ an ideal generated by an $R$-sequence of length at least $2$. Set $d=\operatorname{depth}_{I}(R)$, and let $I$ be an ideal containing $J$ such that $e=\operatorname{depth}_{I}(R/J)\geq 2$. Suppose that $\operatorname{Ass}H_{I}^{d+1}(R)$ is finite and that there is an $R/J$-sequence $x_{1},\ldots,x_{e-1}\in I$ such that $R/(x_{1},\ldots,x_{e-1},J)$ is a locally almost factorial normal ring. Then,
$\operatorname{Ass}H_{I}^{d+1}(J)\text{ is finite}\Leftrightarrow\operatorname{Ass}H_{I}^{d}(R/J)\text{ is finite.}$
In Section 5, we consider the application of the theorem to ideals $I\supset J$, each of which is generated by a part of a regular system of parameters of a regular local ring $(R,\mathfrak{m})$. As a consequence, we obtain the following result, the proof of which is provided in Section 5:
Theorem 1.5.
Let $(R,\mathfrak{m})$ be an LC-finite regular local ring and $I\supset J$ ideals of $R$ with $\operatorname{depth}_{J}(R)\geq 2$. Suppose that each of $I$ and $J$ is generated by a part of a regular system of parameters. Then, for any $i\geq 1$,
$\operatorname{Ass}H_{I}^{i}(J)\text{ is finite}\Leftrightarrow\operatorname{Ass}H_{I}^{i-1}(R/J)\text{ is finite.}$
Over Noetherian local rings $R$ of small dimension, Marley[Mar] has shown several results regarding the finiteness of $\operatorname{Ass}H_{I}^{i}(M)$ for finite $R$-modules $M$. Using Marley’s result, which establishes the finiteness of $\operatorname{Ass}H_{I}^{i-1}(R/J)$, we deduce the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ by applying the theorem above (see Proposition 5.3, Proposition 5.4).
2 Background
2.1 Associated primes
Let $R$ be a ring, and $M$ an $R$-module. A prime ideal $P$ of $R$ is called an associated prime of $M$ if $P=\operatorname{ann}(x)$ for some $x\in M$. The set of associated primes of $M$ is written $\operatorname{Ass}M$, or $\operatorname{Ass}_{R}M$.
Remark 2.1.
Let $A\to B$ be a ring homomorphism, and $M$ a $B$-module. If $A\to B$ is surjective, $B=A/I$ for some ideal $I$ of $A$. $\operatorname{Spec}A/I$ is identified with the Zariski closed subset $V(I)$ of $\operatorname{Spec}A$, and $\operatorname{Ass}_{A/I}M=\operatorname{Ass}_{A}M$ as a subset of $\operatorname{Spec}A$. Let $S$ be a multiplicative set of $A$, and $B=A_{S}$. $\operatorname{Spec}B$ is identified with a subset of $\operatorname{Spec}A$. If A is Noetherian, $\operatorname{Ass}_{A_{S}}M=\operatorname{Ass}_{A}M$ as a subset of $\operatorname{Spec}A$ [Mat, Theorem 6.2].
Let $R$ be a ring, $M$ an $R$-module, and $P\in\operatorname{Ass}M$. Since there exists an injection $A/P\hookrightarrow M$, $0\to A_{P}/PA_{P}\to M_{P}$ is exact, where $A_{P}/PA_{P}\neq 0$. Thus, $M_{P}\neq 0$ and $P\in\operatorname{Supp}M$. This implies $\operatorname{Ass}M\subset\operatorname{Supp}M$.
2.2 Local cohomology modules
Let $R$ be a ring, $I$ an ideal of $R$, and $M$ an $R$-module. Set
$H_{I}^{0}(M)=\{x\in M\mid\text{there exists $n>0$ such that $I^{n}x=0$}\}.$
$H_{I}^{0}$ is a left exact functor. Let $H_{I}^{i}$ be the $i$-th right derived functor of $H_{I}^{0}$. $H_{I}^{i}(M)$ is called the $i$-th local cohomology module of M with respect to I.
From the definition, $H_{I}^{i}(M)=H_{\sqrt{I}}^{i}(M)$ and every element of $H_{I}^{i}(M)$ is annihilated by some power of $I$. If $P\in\operatorname{Spec}R$ does not contain $I$, a multiplicative set $R\setminus P$ contains some element $x\in I$. Since any element of $H_{I}^{i}(M)$ is annihilated by some power of $x$, $H_{I}^{i}(M)_{P}=0$. Therefore, $\operatorname{Supp}H_{I}^{i}(M)\subset V(I)$, and $\operatorname{Ass}H_{I}^{i}(M)\subset V(I)$.
Let $R$ be a Noetherian ring, $I=(x_{1},\ldots,x_{t})$ an ideal of $R$, and $M$ an $R$-module. A local cohomology module $H_{I}^{i}(M)$ is computed as the $i$-th cohomology of the Čech complex, which has nonzero terms only in the range $0\leq i\leq t$, see [Eis, Theorem A1.3]. From this, the following holds:
Proposition 2.2.
Let $R$ be a Noetherian ring, $I=(x_{1},\ldots,x_{t})$ an ideal of $R$, and $M$ an $R$-module. If $i>t$, then $H_{I}^{i}(M)=0$.
For an ideal $I$ of a Noetherian ring $R$, the arithmetic rank of $I$, written $\operatorname{ara}(I)$, is defined by
$\operatorname{ara}(I)=\text{min}\{n\geq 0\mid\text{there exists $x_{1},\ldots,x_{n}$ such that $\sqrt{I}=\sqrt{(x_{1},\ldots,x_{n})}$}\}.$
For the zero ideal, set $\operatorname{ara}(0)$ to be 0. The proposition states that if $R$ is Noetherian, then for any ideal $I$ of $R$ and any $R$-module $M$, $H_{I}^{i}(M)=0$ for all $i>\operatorname{ara}(I)$.
Lemma 2.3.
For ideals $I_{1},\ldots,I_{t}$ of a Noetherian ring $R$,
$\sqrt{I_{1}\cap\cdots\cap I_{t}}=\sqrt{I_{1}\cdots I_{t}}.$
Proof.
By induction on $t$, it suffices to show that $\sqrt{I_{1}\cap I_{2}}=\sqrt{I_{1}I_{2}}$. It is obvious $\sqrt{I_{1}I_{2}}\subset\sqrt{I_{1}\cap I_{2}}$. If $a\in\sqrt{I_{1}\cap I_{2}}$, $a^{n}\in I_{1}\cap I_{2}$ for some $n>0$. Then, $a^{2n}=a^{n}a^{n}\in I_{1}I_{2}$, and $a\in\sqrt{I_{1}I_{2}}$. ∎
The following two results are obtained since a local cohomology module can be computed by using the Čech complex, see [Hun, Proposition 2.14].
Proposition 2.4.
Let $R$ be a Noetherian ring, $I$ an ideal of $R$, and $M$ an $R$-module. If $R\to S$ is a flat ring homomorphism, $H_{I}^{i}(M)\otimes_{R}S=H_{IS}^{i}(M\otimes_{R}S)$.
Proposition 2.5.
Let $R$ be a Noetherian ring, $I$ an ideal of $R$, and $R\to S$ a ring homomorphism. Let $M$ be a $S$-module, then $H_{I}^{i}(M)=H_{IS}^{i}(M)$.
The following theorem exhibits the relationship among dimension, depth and the vanishing of a local cohomology module, see [Hun, Theorem 5.8].
Theorem 2.6.
Let $R$ be a Noetherian ring, $I$ an ideal of $R$, and $M$ a finite $R$-module. Then, $H_{I}^{i}(M)=0$, for $i<\operatorname{depth}_{I}(M)$, and for $i>\dim M$.
2.3 Almost factorial rings
Definition 2.7 ([Mat, Section 12]).
Let $A$ be an integral domain and $K$ its field of fractions. Let $K^{\ast}$ be the multiplicative group of $K$. Let $\mathcal{F}=\{R_{\lambda}\}_{\lambda\in\Lambda}$ be a family of DVRs of $K$ and $v_{\lambda}$ the normalized valuation corresponding to $R_{\lambda}$. If the following two conditions hold, $A$ is called a Krull ring, and $\mathcal{F}$ is called a defining family of $A$ :
  1. 1.
    $A=\underset{\lambda\in\Lambda}{\bigcap}R_{\lambda}$.
  2. 2.
    for every $x\in K^{\ast}$, there are at most a finite number of $\lambda\in\Lambda$ such that $v_{\lambda}(x)\neq 0$.
A Noetherian normal domain is known to be a Krull ring. Let $\mathcal{F}$ be a defining family of a Krull ring $A$, and $P$ a height 1 prime ideal of $A$. Then, $A_{P}\in\mathcal{F}$ (see [Mat, Theorem 12.3]), and $A_{P}$ is a DVR. Let $\mathcal{P}$ be the set of height 1 prime ideals of $A$. Let $D(A)$ be the free Abelian group generated by $\mathcal{P}$, that is, every element of $D(A)$ is written as a formal sum
$\sum_{P\in\mathcal{P}}n_{P}\cdot P,\quad n_{P}\in\mathbb{Z},$
where all but finitely many $n_{P}=0$ and addition is defined by
$\sum_{P\in\mathcal{P}}n_{P}\cdot P+\sum_{P\in\mathcal{P}}n^{\prime}_{P}\cdot P=\sum_{P\in\mathcal{P}}(n_{P}+n^{\prime}_{P})\cdot P.$
Let $K$ be the field of fractions of $A$ and $K^{\ast}$ the multiplicative group of $K$. For $a\in K^{\ast}$, set $\operatorname{div}(a)=\sum_{P\in\mathcal{P}}v_{P}(a)\cdot P$, where $v_{P}$ is the normalized valuation corresponding to $A_{P}$. div : $K^{\ast}\to D(A)$ is a group homomorphism. The divisor class group of $A$, written $C(A)$, is defined by $C(A)=D(A)/\operatorname{div}(K^{\ast})$. In fact, $A$ is a UFD if and only if C(A)C(A)=0.
Definition 2.8.
Krull ring $A$ is said to be almost factorial if $C(A)$ is a torsion group, that is, every element of $C(A)$ has finite order.
Note that UFDs are almost factorial. The following property of almost factorial rings enables us to rewrite local cohomology modules into more tractable form, which facilitates their computation (see Lemma 3.5, Proposition 3.6).
Proposition 2.9.
Let $A$ be an almost factorial Krull ring. If $I$ is an ideal of pure hight $1$, that is, every minimal prime ideal of $I$ has height $1$, then $\sqrt{I}=\sqrt{aA}$ for some $a\in A$.
Proof.
Let $P_{1},\ldots,P_{t}$ be all the minimal prime ideals of $I$, then $\sqrt{I}=P_{1}\cap\cdots\cap P_{t}$. For an each $P_{i}$, we show that $P_{i}=\sqrt{a_{i}A}$ for some $a_{i}\in A$. Since $I$ is of pure hight 1, $P_{i}$ is a hight 1 prime ideal of $A$. Let $K$ be the field of fractions of $A$ and $K^{\ast}$ the multiplicative group of $K$. Since $C(A)$ is torsion, $n\cdot P_{i}\in\operatorname{div}(K^{\ast})$ for some $n\in\mathbb{N}_{>0}$, that is, for some $a_{i}\in K^{\ast}$,
$n\cdot P_{i}=\operatorname{div}(a_{i})=\sum_{P\in\mathcal{P}}v_{P}(a_{i})\cdot P.$
It means that for $P\in\mathcal{P}$,
$v_{P}(a_{i})=\begin{cases}n&\text{if $P=P_{i}$}\\ 0&\text{if $P\neq P_{i}$}\end{cases}.$
Then, $a_{i}\in P_{i}A_{P_{i}}$ and $a_{i}\in A_{P}\setminus PA_{P}$ if $P\neq P_{i}$. In particular, $a_{i}\in\bigcap_{P\in\mathcal{P}}A_{P}$. Since the family $\{A_{P}\mid P\in\mathcal{P}\}$ of DVRs is a defining family of $A$, $a_{i}\in P_{i}A_{P_{i}}\cap A=P_{i}$. Thus, $\sqrt{a_{i}A}\subset P_{i}$. Furthermore,
$a_{i}A_{P}=\begin{cases}P_{i}^{n}A_{P_{i}}&\text{if $P=P_{i}$}\\ A_{P}&\text{if $P\neq P_{i}$}\end{cases}.$
It follows that
$a_{i}A=\bigcap_{P\in\mathcal{P}}(a_{i}A_{P}\cap A)=P_{i}^{n}A_{P_{i}}\cap A.$
If $x\in P_{i}$, then $x^{n}\in a_{i}A$. Hence, $P_{i}=\sqrt{a_{i}A}$. By Lemma 2.3,
$\sqrt{I}=\sqrt{a_{1}A}\cap\ldots\cap\sqrt{a_{t}A}=\sqrt{a_{1}\cdots a_{t}A}.$
3 Theoretical tools
In this section, we briefly summarize the Lewis’s approach to Question 1.2, which forms the basis for our discussion. We restate the question as follows: Let $R$ be an LC-finite regular ring, $J$ an ideal of $R$ generated by an $R$-sequence, and $I\supset J$. Fix $i\geq 1$. Then, $\operatorname{Ass}H_{I}^{i}(J)$ is finite if $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite?
Since both $\operatorname{Ass}H_{I}^{0}(R/J)$ and $\operatorname{Ass}H_{I}^{1}(J)$ are finite, the question is true when $i=1$. Let $j>0$ be the length of an $R$-sequence generating $J$. In the case $j=1$, as an $R$-module, $J\simeq R$ and $H_{I}^{i}(J)\simeq H_{I}^{i}(R)$. The LC-finiteness of $R$ implies that $\operatorname{Ass}H_{I}^{i}(J)$ is finite, and the claim is trivial. We may assume $j\geq 2$.
For a Noetherian normal ring $R$, a local ring $R_{P}$ at $P\in\operatorname{Spec}R$ is a Noetherian normal domain, hence a Krull ring. If $R_{P}$ is almost factorial for any $P\in\operatorname{Spec}R$, $R$ is said to be locally almost factorial.
Theorem 3.1 ([Lew, Theorem 2.1]).
Let $R$ be a locally almost factorial Noetherian normal ring, and $I$, $J$ be ideals of $R$. The set $\operatorname{Ass}H_{I}^{2}(J)$ is finite.
Since a regular ring is locally almost factorial, Theorem 3.1 implies that $\operatorname{Ass}H_{I}^{2}(J)$ is always finite in the above question. Since $\operatorname{Ass}H_{I}^{1}(R/J)$ is finite, the question is true when $i=2$. The assertions for $i=3,4$ also hold under the hypothesis that $R/J$ is normal and locally almost factorial [Lew, Theorem 4.2(iii)]. The proof begins by reducing the problem to the case $\operatorname{depth}_{I}(R)\geq i-1$, using a generalized version of Hellus’s isomorphism (see [Lew, Theorem 4.1 and Corollary 4.3]). Furthermore,
Lemma 3.2 ([Lew, Lemma 4.1]).
Let $R$ be a Noetherian ring and $I$, $J$ be ideals of $R$. Fix $i\geq 1$ and assume $\operatorname{Ass}H_{I}^{i}(R)$ is finite. If $\operatorname{depth}_{I}(R)>i-1$, then $\operatorname{Ass}H_{I}^{i}(J)$ is finite if and only if $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite.
By the lemma, we should view only the case $\operatorname{depth}_{I}(R)=i-1$. Since
$i-1=\operatorname{depth}_{I}(R)=\operatorname{depth}_{I}(R/J)+\operatorname{depth}_{J}(R),$
and $j=\operatorname{depth}_{J}(R)\geq 2$, there are the only cases
$(\operatorname{depth}_{I}(R/J),\operatorname{depth}_{J}(R))=\begin{cases}(0,2)&\text{for $i=3$}\\ (1,2),(0,3)&\text{for $i=4$}\end{cases}.$
For the case $\operatorname{depth}_{I}(R/J)=0$ and the case $\operatorname{depth}_{I}(R/J)=1$, the following results are found in the proof of [Lew, Theorem 4.2]:
Theorem 3.3 (cf. [Lew, Theorem 4.2]).
Let $R$ be a Noetherian ring, $J$ an ideal generated by an $R$-sequence of length at least $2$, and $I$ an ideal containing $J$. Set $d=\operatorname{depth}_{I}(R)$. Suppose that $\operatorname{Ass}H_{I}^{d+1}(R)$ is finite. Then, in the following two cases, $\operatorname{Ass}H_{I}^{d+1}(J)$ is finite if and only if $\operatorname{Ass}H_{I}^{d}(R/J)$ is finite.
  1. 1.
    In the case $\operatorname{depth}_{I}(R/J)=0$, the claim is true if $R/J$ is Cohen-Macaulay and a locally domain.
  2. 2.
    In the case $\operatorname{depth}_{I}(R/J)=1$, the claim is true if $R/J$ is a locally almost factorial normal ring.
Later, we generalize Case 2 by following Lewis’s approach in [Lew] and incorporating the method to remove the restriction on $\operatorname{depth}_{I}(R/J)$. For the sake of discussion, the proof of Theorem 3.3 is given below. First, we prepare preliminary results.
Lemma 3.4.
Let $A$ be a Cohen-Macaulay ring, $I$ an ideal of $A$ with $\operatorname{depth}_{I}(A)=0$. If $A$ is a locally domain, $\sqrt{I}=\sqrt{aA}\cap I_{0}$ for some $a\in A$ and for some ideal $I_{0}$ with $\operatorname{depth}_{I_{0}}(A)>0$.
Proof.
Since $\operatorname{ht}I=\operatorname{depth}_{I}(A)=0$, we can write $\sqrt{I}=J\cap I_{0}$, in which $J$ is a finite intersection of minimal prime ideals of $A$, and $I_{0}$ is a finite intersection of height >0>0 prime ideals. Thus, $\operatorname{depth}_{I_{0}}(A)=\operatorname{ht}I_{0}>0$. To see that $J$ is a radical of some principal ideal of $A$, we show that every minimal prime ideal of $A$ is a radical of some principal ideal of $A$. Let $P_{1},\ldots,P_{t}$ be all the minimal prime ideals of $A$. Since $A$ is Noetherian and a locally domain,
$A\simeq A/P_{1}\times\cdots\times A/P_{t}.$
Let $e_{i}\in A$ be the element corresponding to “the $i$-th standard idempotent” $(0,\ldots,0,1,0,\ldots,0)\in A/P_{1}\times\cdots\times A/P_{t}$. Each $P_{i}$ corresponds to the ideal
$A/P_{1}\times\cdots\times A/P_{i-1}\times(0)\times A/P_{i+1}\times\cdots\times A/P_{t}.$
Thus, $e_{i}\notin P_{i}$, and $e_{j}\in P_{i}$ if $j\neq i$. In addition, $e_{1}+\cdots+e_{t}=1$, since this element corresponds to $(1,\ldots,1)$. Hence, $1-e_{i}\in P_{i}$, and it follows that $P_{i}$ is the only minimal element of $V((1-e_{i})A)$. Indeed, if not, there exists $P\in V((1-e_{i})A)$ such that $P\supset P_{j}$ for some $j\neq i$. Then, $1-e_{i},e_{i}\in P$, and $1\in P$, yielding a contradiction. ∎
Lemma 3.5 ([Lew, Corollary 2.1]).
Let $A$ be a locally almost factorial Noetherian normal ring, and $I$ an ideal of $A$ such that $\operatorname{depth}_{I}(A)=1$. Then, there is an ideal $I_{0}$ of $A$ such that
$\operatorname{depth}_{I_{0}}(A)>1,$
and a finite open covering $\{\operatorname{Spec}(A_{f_{1}}),\ldots,\operatorname{Spec}(A_{f_{t}})\mid f_{1},\ldots,f_{t}\in A\}$ of $\operatorname{Spec}A$ such that for each $i=1,\ldots,t$,
$\sqrt{IA_{f_{i}}}=\sqrt{a_{i}A_{f_{i}}}\cap I_{0}A_{f_{i}}\quad\text{for some $a_{i}\in A$}.$
Proposition 3.6 ([Lew, Corollary 2.2]).
Let $A$ be a Noetherian ring, $a\in A$, and $I_{0}$ an ideal of $A$. Let $I=\sqrt{aA}\cap I_{0}$. There is a natural isomorphism of functors $H_{I}^{i}(-)\simeq H_{I_{0}}^{i}(-)_{a}$ for all $i\geq 2$.
Lemma 3.7.
Let $\psi:M\to N$ be a homomorphism of $A$-modules, and $B$ an $A$-algebra. If $N$ also has a $B$-module structure, then $\psi$ factors through the natural map $M\to B\otimes_{A}M$, $x\mapsto 1\otimes x$:
$M\xrightarrow{\psi}N\quad=\quad M\to B\otimes_{A}M\to N.$
Proof.
By Hom-Tensor adjointness,
$\operatorname{Hom}_{B}(B\otimes_{A}M,N)\simeq\operatorname{Hom}_{A}(M,\operatorname{Hom}_{B}(B,N))\simeq\operatorname{Hom}_{A}(M,N).$
Under the first isomorphism, $\tilde{f}:B\otimes_{A}M\to N$ corresponds to the map
$F:M\to\operatorname{Hom}_{B}(B,N),\quad x\mapsto(F_{x}\colon B\ni b\mapsto\tilde{f}(b\otimes x)).$
The second isomorphism is induced by
$\operatorname{Hom}_{B}(B,N)\simeq N,\quad g\mapsto g(1).$
Under the composition of these isomorphisms, $\tilde{f}\colon B\otimes_{A}M\to N$ corresponds to the map $f\colon M\to N,x\mapsto\tilde{f}(1\otimes x)$. Hence, there is a map $\tilde{\psi}\colon B\otimes_{A}M\to N$ such that
$M\xrightarrow{\psi}N\quad=\quad M\to B\otimes_{A}M\xrightarrow{\tilde{\psi}}N.$
Proof of Theorem 3.3.
Set $S=R/J$. In Case 1, by Lemma 3.4,
$\sqrt{I/J}=\sqrt{a(R/J)}\cap(I_{0}/J)$
for some $a\in R$ and for some ideal $I_{0}\supset J$ of $R$ with $\operatorname{depth}_{I_{0}}(R/J)>0$. Then, $\sqrt{IS}=\sqrt{aS\cap I_{0}S}$ and $H_{I}^{i}(S)=H_{aS\cap I_{0}S}^{i}(S)$. By Proposition 3.6, $H_{I}^{i}(S)\simeq H_{I_{0}}^{i}(S)_{a}$ for all $i\geq 2$. Now,
$d=\operatorname{depth}_{I}(R)=\operatorname{depth}_{I}(R/J)+\operatorname{depth}_{J}(R)=0+\operatorname{depth}_{J}(R)\geq 2,$
and $H_{I}^{d}(S)$ has a $S_{a}$-module structure. The natural map $\psi\colon H_{I}^{d}(R)\to H_{I}^{d}(S)$ factors through $S_{a}\otimes_{R}H_{I}^{d}(R)$ by Lemma 3.7. We want to show $\psi=0$. It suffices to show that $S_{a}\otimes_{R}H_{I}^{d}(R)=0$. Since
$S_{a}\otimes_{R}H_{I}^{d}(R)\simeq S_{a}\otimes_{R_{a}}R_{a}\otimes_{R}H_{I}^{d}(R)\simeq S_{a}\otimes_{R_{a}}H_{IR_{a}}^{d}(R_{a}),$
it suffices to show $H_{IR_{a}}^{d}(R_{a})=0$. The following two equalities hold:
$\sqrt{I}=\sqrt{(aR+J)\cap I_{0}},\quad\sqrt{IR_{a}}=\sqrt{(aR+J)R_{a}\cap I_{0}R_{a}}=\sqrt{I_{0}R_{a}}.$
By the second equality, $H_{IR_{a}}^{d}(R_{a})\simeq H_{I_{0}R_{a}}^{d}(R_{a})$. Now $J$ is generated by an $R$-sequence and contained in $I_{0}$,
$\operatorname{depth}_{I_{0}R_{a}}(R_{a})\geq\operatorname{depth}_{I_{0}}(R)=\operatorname{depth}_{I_{0}}(R/J)+\operatorname{depth}_{J}(R)>d.$
Therefore, $H_{IR_{a}}^{d}(R_{a})=0$ and $\psi=0$. From the long exact sequence of a local cohomology,
$0\to H_{I}^{d}(S)\to H_{I}^{d+1}(J)\to H_{I}^{d+1}(R)$
is exact, and
$\operatorname{Ass}H_{I}^{d}(S)\subset\operatorname{Ass}H_{I}^{d+1}(J)\subset\operatorname{Ass}H_{I}^{d}(S)\cup\operatorname{Ass}H_{I}^{d+1}(R).$
By the assumption $\operatorname{Ass}H_{I}^{d+1}(R)$ is finite, the claim follows.
In Case 2, by Lemma 3.5, there is an ideal $I_{0}$ of $R$,
$I_{0}\supset J\quad\text{with $\operatorname{depth}_{I_{0}}(R/J)>1$},$
and a finite open covering of $\operatorname{Spec}S$,
$\{\operatorname{Spec}(S)\cap D(f_{1}),\ldots,\operatorname{Spec}(S)\cap D(f_{t})\mid f_{1},\ldots,f_{t}\in R\}$
such that for each $i$,
$\sqrt{IR_{f_{i}}/JR_{f_{i}}}=\sqrt{a_{i}(R_{f_{i}}/JR_{f_{i}})}\cap(I_{0}R_{f_{i}}/JR_{f_{i}})\quad\text{for some $a_{i}\in R$},$
where $D(f_{i})=\operatorname{Spec}R_{f_{i}}$ is the principal open subset of $\operatorname{Spec}R$. For any $R$-module $M$, $\operatorname{Ass}H_{I}^{j}(M)\subset V(I)\subset V(J)=\operatorname{Spec}S$. In addition, for each $i$,
$\operatorname{Ass}H_{I}^{j}(M_{f_{i}})=\operatorname{Ass}H_{I}^{j}(M)\cap D(f_{i}).$
It follows that $\operatorname{Ass}H_{I}^{j}(M)$ is finite if and only if $\operatorname{Ass}H_{I}^{j}(M_{f_{i}})$ is finite for all $i=1,\ldots,t$. Then, we should prove that for all $i=1,\ldots,t$,
$\operatorname{Ass}H_{I}^{d}(S_{f_{i}})\text{ is finite}\Leftrightarrow\operatorname{Ass}H_{I}^{d+1}(JR_{f_{i}})\text{ is finite}.$
Now, $d=1+\operatorname{depth}_{J}(R)\geq 3$. Replacing $R$ by $R_{f_{i}}$ and $a$ by $a_{i}$ in the discussion of Case 1, we see that it suffices to show that $H_{I(R_{f_{i}})_{a_{i}}}^{d}((R_{f_{i}})_{a_{i}})=0$, and we can also conclude $\sqrt{I(R_{f_{i}})_{a_{i}}}=\sqrt{I_{0}(R_{f_{i}})_{a_{i}}}$. Then,
$H_{I(R_{f_{i}})_{a_{i}}}^{d}((R_{f_{i}})_{a_{i}})\simeq H_{I_{0}(R_{f_{i}})_{a_{i}}}^{d}((R_{f_{i}})_{a_{i}})=H_{I_{0}}^{d}((R_{f_{i}})_{a_{i}}).$
It follows that $H_{I(R_{f_{i}})_{a_{i}}}^{d}((R_{f_{i}})_{a_{i}})=0$, since
$\operatorname{depth}_{I_{0}}((R_{f_{i}})_{a_{i}})\geq\operatorname{depth}_{I_{0}}(R)=\operatorname{depth}_{I_{0}}(R/J)+\operatorname{depth}_{J}(R)>d.$
Theorem 3.8 ([Lew, Theorem 4.2]).
Let $R$ be an LC-finite regular ring, $J$ an ideal generated by an $R$-sequence of length at least $2$. Let $I$ be an ideal of $R$ containing $J$. Suppose that $R/J$ is normal and locally almost factorial. Then, for any $i\leq 4$, $\operatorname{Ass}H_{I}^{i}(J)$ is finite if and only if $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite.
Proof.
As we have already seen at the beginning of this section, we only need to consider the following cases, where $\operatorname{depth}_{I}(R/J)=0,1$:
$(\operatorname{depth}_{I}(R/J),\operatorname{depth}_{J}(R))=(0,2),(0,3),(1,2).$
$R/J$ is Cohen-Macaulay, since $R$ is regular and $J$ is generated by an $R$-sequence. In addition, by the assumption that $R/J$ is normal, $R/J$ is a locally domain. For all of the cases above, the claim holds by Theorem 3.3. ∎
4 Main results
In this section, we consider the case $\operatorname{depth}_{I}(R/J)\geq 2$. To handle this case in accordance with Lewis’s approach, we assume the existence of a suitable $R/J$-sequence in $I$. The assumption in the following statement coincides the assumption in Case 2 of Theorem 3.3, when $\operatorname{depth}_{I}(R/J)=1$:
Proposition 4.1.
Let $R$ be a Noetherian ring, $J$ an ideal generated by an $R$-sequence of length at least $2$. Set $S=R/J$. Let $I$ be an ideal containing $J$ such that $e=\operatorname{depth}_{I}(R/J)\geq 2$. Suppose that there is an $R/J$-sequence $x_{1},\ldots,x_{e-1}\in I$ such that $R/(x_{1},\ldots,x_{e-1},J)$ is a locally almost factorial normal ring. Then, there is an ideal $I_{0}\supset(x_{1},\ldots,x_{e-1},J)$ of $R$ such that
$\operatorname{depth}_{I_{0}}(S)>e$
and a finite open covering $\{V(I)\cap D(f_{1}),\ldots,\ V(I)\cap D(f_{t})\mid f_{1},\ldots,f_{t}\in R\}$ of $V(I)$ such that for each $i=1,\ldots,t$
$\sqrt{IS_{f_{i}}}=\sqrt{(x_{1},\ldots,x_{e-1},y_{i})S_{f_{i}}}\cap I_{0}S_{f_{i}}=\sqrt{y_{i}S_{f_{i}}\cap I_{0}S_{f_{i}}+(x_{1},\ldots,x_{e-1})S_{f_{i}}}$
for some $y_{i}\in R$. Where, $D(f_{i})$ is the principal open subset of $\operatorname{Spec}R$.
Proof.
With the notation $\underline{x}=(x_{1},\ldots,x_{e-1})$, the following equality holds:
$\operatorname{depth}_{I}(R/(\underline{x},J))=\operatorname{depth}_{I}(R/J)-(e-1)=1$
By Lemma 3.5, there is an ideal $I_{0}\supset(\underline{x},J)$ of $R$ such that $\operatorname{depth}_{I_{0}}(R/(\underline{x},J))>1$, and a finite open covering of $\operatorname{Spec}R/(\underline{x},J)$
$\{\operatorname{Spec}(R/(\underline{x},J))\cap D(f_{1}),\ldots,\operatorname{Spec}(R/(\underline{x},J))\cap D(f_{t})\mid f_{1},\ldots,f_{t}\in R\}$
such that for each $i=1,\ldots,t$
$\sqrt{I(R/(\underline{x},J))_{f_{i}}}=\sqrt{y_{i}(R/(\underline{x},J))_{f_{i}}}\cap I_{0}(R/(\underline{x},J))_{f_{i}}\quad\text{for some $y_{i}\in R$}.$
Then, $\operatorname{depth}_{I_{0}}(S)=(e-1)+\operatorname{depth}_{I_{0}}(R/(\underline{x},J))>e$. The above formula is rewritten in the following form: $\sqrt{IS_{f_{i}}/\underline{x}S_{f_{i}}}=\sqrt{y_{i}(S_{f_{i}}/\underline{x}S_{f_{i}})}\cap(I_{0}S_{f_{i}}/\underline{x}S_{f_{i}})$. From this, we derive the first equality in the claim. Set $S_{i}=S_{f_{i}}$. Let $a\in\sqrt{IS_{i}}$, then, $a\mod\underline{x}S_{i}\in\sqrt{IS_{f_{i}}/\underline{x}S_{f_{i}}}=\sqrt{y_{i}(S_{i}/\underline{x}S_{i})}\cap(I_{0}S_{i}/\underline{x}S_{i})$. It follows that $a\in I_{0}S_{i}$, and $a^{n}-y_{i}b\in\underline{x}S_{i}$ for some $n>0$ and $b\in S_{i}$. Thus, $a\in\sqrt{(\underline{x},y_{i})S_{i}}\cap I_{0}S_{i}$. Conversely, let $a\in\sqrt{(\underline{x},y_{i})S_{i}}\cap I_{0}S_{i}$. Then, $a\mod\underline{x}S_{i}\in\sqrt{y_{i}(S_{i}/\underline{x}S_{i})}\cap(I_{0}S_{i}/\underline{x}S_{i})=\sqrt{IS_{i}/\underline{x}S_{i}}$, and $a^{n}\in IS_{i}+\underline{x}S_{i}$ for some $n>0$. Since $(x_{1},\ldots,x_{e-1})\subset I$, $a\in\sqrt{IS_{i}}$. Therefore, $\sqrt{IS_{i}}=\sqrt{(\underline{x},y_{i})S_{i}}\cap I_{0}S_{i}$.
Next, we prove the second equality in the claim, which, with the notation $S_{i}=S_{f_{i}}$, takes the following form:
$\sqrt{(\underline{x},y_{i})S_{i}}\cap I_{0}S_{i}=\sqrt{y_{i}S_{i}\cap I_{0}S_{i}+\underline{x}S_{i}}.$
Taking the radical of both ideals, it suffices to show that
$\sqrt{(\underline{x},y_{i})S_{i}\cap I_{0}S_{i}}=\sqrt{y_{i}S_{i}\cap I_{0}S_{i}+\underline{x}S_{i}}.$
For any $a$ in the left-hand side, $a^{n}=x_{1}b_{1}+\cdots+x_{e-1}b_{e-1}+y_{i}c\in I_{0}S_{i}$ for some $n>0$ and $b_{1},\ldots,b_{e-1},c\in S_{i}$. Here, $y_{i}c\in y_{i}S_{i}\cap I_{0}S_{i}$, since $x_{1},\ldots x_{e-1}\in I_{0}$. Then, $a^{n}\in y_{i}S_{i}\cap I_{0}S_{i}+\underline{x}S_{i}$. Conversely, for any $a$ in the right-hand side, $a^{n}=x_{1}b_{1}+\cdots+x_{e-1}b_{e-1}+y_{i}c$ for some $n>0$, where $b_{1},\ldots,b_{e-1},c\in S_{i}$, and $y_{i}c\in I_{0}S_{i}$. Then, $a^{n}\in(\underline{x},y_{i})S_{i}\cap I_{0}S_{i}$, since $x_{1},\ldots,x_{e-1}\in I_{0}$.
From the above, we have shown both the inequality $\operatorname{depth}_{I_{0}}(S)>e$ and the equality within the claim. Furthermore, since $\operatorname{Spec}R/(\underline{x},J)\supset V(I)$,
$\{V(I)\cap D(f_{1}),\ldots,\ V(I)\cap D(f_{t})\mid f_{1},\ldots,f_{t}\in R\}$
is an open covering of $V(I)$, as desired. ∎
The relationship of the ideal’s radical given in the proposition shows, through the form of the equation $\sqrt{IS_{i}}=\sqrt{y_{i}S_{i}\cap I_{0}S_{i}+(x_{1},\ldots,x_{e-1})S_{i}}$, how far the structure of $IS_{i}$ is from that of $y_{i}S_{i}\cap I_{0}S_{i}$. In Lewis’s approach, the ideal $y_{i}S_{i}\cap I_{0}S_{i}$ plays an important role in establishing the equivalence between the finiteness of $H_{I}^{d+1}(J)$ and that of $H_{I}^{d}(R/J)$. The difference between $IS_{i}$ and $y_{i}S_{i}\cap I_{0}S_{i}$, namely the $(x_{1},\ldots,x_{e-1})S_{i}$ part, can be ignored for the purpose of computing the local cohomology modules relevant to the present discussion:
Lemma 4.2.
Let $A$ be a Noetherian ring, and $I_{0}$ an ideal of $A$. An ideal $I$ of $R$ is assumed to be of the form
$\sqrt{I}=\sqrt{x_{e}A\cap I_{0}+(x_{1},\ldots,x_{e-1})A}$
for some $x_{1},\ldots,x_{e}\in A$ such that $(x_{e}x_{1},x_{e}x_{2},\ldots,x_{e}x_{e-1})A\subset I_{0}$. Then, for any $A$-module $M$ and any $i>e$, $H_{I}^{i}(M)\simeq H_{x_{e}A\cap I_{0}}^{i}(M)$.
Proof.
By the assumption and with the notation $\underline{x}=(x_{1},\ldots,x_{e-1})$, we have $H_{I}^{i}(M)\simeq H_{x_{e}A\cap I_{0}+\underline{x}A}^{i}(M)$. It follows from the Mayor-Vietoris sequence that
$H_{x_{e}A\cap I_{0}\cap\underline{x}A}^{i-1}(M)\to H_{I}^{i}(M)\to H_{x_{e}A\cap I_{0}}^{i}(M)\oplus H_{\underline{x}A}^{i}(M)\to H_{x_{e}A\cap I_{0}\cap\underline{x}A}^{i}(M)$
is exact. By the assumption that $(x_{e}x_{1},\ldots,x_{e}x_{e-1})A\subset I_{0}$,
$\begin{split}\sqrt{x_{e}A\cap I_{0}\cap\underline{x}A}&=\sqrt{I_{0}\cap\sqrt{x_{e}A\cap\underline{x}A}}=\sqrt{I_{0}\cap\sqrt{(x_{e}x_{1},\ldots,x_{e}x_{e-1})A}}\\ &=\sqrt{I_{0}\cap(x_{e}x_{1},\ldots,x_{e}x_{e-1})A}=\sqrt{(x_{e}x_{1},\ldots,x_{e}x_{e-1})A}\end{split}$
The above exact sequence is rewritten in the following form:
$H_{(x_{e}\underline{x})}^{i-1}(M)\to H_{I}^{i}(M)\to H_{x_{e}A\cap I_{0}}^{i}(M)\oplus H_{\underline{x}A}^{i}(M)\to H_{(x_{e}\underline{x})}^{i}(M),$
where $(x_{e}\underline{x})$ denotes the ideal $(x_{e}x_{1},\ldots,x_{e}x_{e-1})A$. If $i>e$, by Proposition 2.2, this exact sequence is
$0\to H_{I}^{i}(M)\to H_{x_{e}A\cap I_{0}}^{i}(M)\oplus 0\to 0.$
Hence, $H_{I}^{i}(M)\simeq H_{x_{e}A\cap I_{0}}^{i}(M)$ if $i>e$. ∎
At this point, we generalize Case 2 of Theorem 3.3.
Theorem 4.3.
Let $R$ be a Noetherian ring, $J$ an ideal generated by an $R$-sequence of length at least $2$. Set $d=\operatorname{depth}_{I}(R)$, and let $I$ be an ideal containing $J$ such that $e=\operatorname{depth}_{I}(R/J)\geq 2$. Suppose that $\operatorname{Ass}H_{I}^{d+1}(R)$ is finite, and that there is an $R/J$-sequence $x_{1},\ldots,x_{e-1}\in I$ such that $R/(x_{1},\ldots,x_{e-1},J)$ is a locally almost factorial normal ring. Then,
$\operatorname{Ass}H_{I}^{d+1}(J)\text{ is finite}\Leftrightarrow\operatorname{Ass}H_{I}^{d}(R/J)\text{ is finite.}$
Proof.
By Proposition 4.1 and, with the notation $\underline{x}=(x_{1},\ldots,x_{e-1})$ and $S=R/J$, there is an ideal $I_{0}\supset(x_{1},\ldots,x_{e-1},J)$ of $R$ such that
$\operatorname{depth}_{I_{0}}(S)>e,$
and a finite open covering $\{V(I)\cap D(f_{1}),\ldots,V(I)\cap D(f_{t})\mid f_{1},\ldots,f_{t}\in R\}$ of $V(I)$ such that for each $i=1,\ldots,t$
$\sqrt{IS_{f_{i}}}=\sqrt{(\underline{x},y_{i})S_{f_{i}}}\cap I_{0}S_{f_{i}}=\sqrt{y_{i}S_{f_{i}}\cap I_{0}S_{f_{i}}+\underline{x}S_{f_{i}}}$
for some $y_{i}\in R$. By Lemma 4.2,
$H_{I}^{j}({S_{f_{i}}})\simeq H_{y_{i}R\cap I_{0}}^{j}(S_{f_{i}})\quad\text{for all $j>e$}.$
Since both $\operatorname{Ass}H_{I}^{d}(S)$ and $\operatorname{Ass}H_{I}^{d+1}(J)$ are contained in $V(I)$, we should prove that for $i=1,\ldots,t$,
$\operatorname{Ass}H_{I}^{d}(S_{f_{i}})\text{ is finite}\Leftrightarrow\operatorname{Ass}H_{I}^{d+1}(JR_{f_{i}})\text{ is finite}.$
By Proposition 3.6, $H_{I}^{j}({S_{f_{i}}})\simeq H_{I_{0}}^{j}(S_{f_{i}})_{y_{i}}$ for all $j>e$ $(\geq 2)$. $H_{I}^{d}(S_{f_{i}})$ has a $(S_{f_{i}})_{y_{i}}$-module structure, since $d=e+\operatorname{depth}_{J}(R)\geq e+2$. As we have seen in the proof of Theorem 3.3, it suffices to show that $H_{I}^{d}((R_{f_{i}})_{y_{i}})=0$ and we can also conclude
$\sqrt{I(R_{f_{i}})_{y_{i}}}=\sqrt{(\underline{x},y_{i},J)(R_{f_{i}})_{y_{i}}\cap I_{0}(R_{f_{i}})_{y_{i}}}=\sqrt{I_{0}(R_{f_{i}})_{y_{i}}}.$
It follows from this formula that $H_{I}^{d}((R_{f_{i}})_{y_{i}})\simeq H_{I_{0}}^{d}((R_{f_{i}})_{y_{i}})=0$, since
$\operatorname{depth}_{I_{0}}((R_{f_{i}})_{y_{i}})\geq\operatorname{depth}_{I_{0}}(R)=\operatorname{depth}_{I_{0}}(S)+\operatorname{depth}_{J}(R)>d.$
Corollary 4.4.
Let $R$ be a Noetherian ring, $J$ an ideal generated by $R$-sequence of length at least $2$. Set $d=\operatorname{depth}_{I}(R)$, and let $I$ be an ideal containing $J$ such that $e=\operatorname{depth}_{I}(R/J)\geq 2$. Suppose that $\operatorname{Ass}H_{I}^{d}(R)$, $\operatorname{Ass}H_{I}^{d+1}(R)$ are finite, and that there is an $R/J$-sequence $x_{1},\ldots,x_{e-1}\in I$ such that $R/(x_{1},\ldots,x_{e-1},J)$ is a locally almost factorial normal ring. Then, for all $i\leq d+1$,
$\operatorname{Ass}H_{I}^{i}(J)\text{ is finite}\Leftrightarrow\operatorname{Ass}H_{I}^{i-1}(R/J)\text{ is finite.}$
Proof.
For the case $i=d+1$, it follows by Theorem 4.3. If $i\leq d$,
$0\to H_{I}^{i-1}(R/J)\to H_{I}^{i}(J)\to H_{I}^{i}(R)$
is exact. If $i<d$, $H_{I}^{i}(R)=0$, and the claim immediately follows. If $i=d$, the claim follows from the assumption that $\operatorname{Ass}H_{I}^{d}(R)$ is finite. ∎
5 Applications
Let $(R,\mathfrak{m})$ be an $r$-dimensional regular local ring. $x_{1},\ldots,x_{r}\in\mathfrak{m}$ is a regular system of parameters if $\mathfrak{m}=(x_{1},\ldots,x_{r})$, or equivalently if the images of $x_{1},\ldots,x_{r}$ in $\mathfrak{m}/\mathfrak{m}^{2}$ are linearly independent over $R/\mathfrak{m}$. For any ideal $I$ of $R$, $R/I$ is regular if and only if $I$ is generated by a part of a regular system of parameters. Let $I\supset J$ be ideals of $R$ with $\operatorname{depth}_{J}(R)\geq 2$. When each of $I$ and $J$ is generated by a part of a regular system of parameters, we can take an $R/J$-sequence as stated in the claim of Theorem 4.3:
Lemma 5.1.
Let $(R,\mathfrak{m},k)$ be a regular local ring and $I\supset J$ ideals of $R$. If each of $I$ and $J$ is generated by a part of a regular system of parameters, there exists a subset of a regular system of parameters $x_{1},\ldots,x_{j},\ldots,x_{d}\in\mathfrak{m}$ such that $J=(x_{1}\ldots,x_{j})$ and $I=(x_{1},\ldots,x_{j},\ldots,x_{d})$.
Proof.
Since $(R/J)/(I/J)\simeq R/I$ is regular, $I/J$ is generated by a part of a regular system of parameters of $R/J$. We can take $y_{1},\ldots,y_{n}\in I$ such that $I/J=(y_{1},\ldots,y_{n})R/J$, and the images of $y_{1},\ldots,y_{n}$ in $(\mathfrak{m}/J)/(\mathfrak{m}/J)^{2}$ are linearly independent over $k$. Since $\mathfrak{m}/(\mathfrak{m}^{2}+J)\simeq(\mathfrak{m}/J)/(\mathfrak{m}/J)^{2}$ as a $k$-module, the images of $y_{1},\ldots,y_{n}$ in $\mathfrak{m}/(\mathfrak{m}^{2}+J)$ are also linearly independent over $k$. Let $x_{1},\ldots,x_{j}\in\mathfrak{m}$ be a part of a regular system of parameters which generates $J$. Then, $I=(x_{1},\ldots,x_{j},y_{1},\ldots,y_{n})$. It suffices to show that the images of $x_{1},\ldots,x_{j},y_{1},\ldots,y_{n}$ in $\mathfrak{m}/\mathfrak{m}^{2}$ are linearly independent over $k$. Assume that $a_{1},\ldots,a_{j},b_{1},\ldots,b_{n}\in R$, and
$a_{1}x_{1}+\cdots+a_{j}x_{j}+b_{1}y_{1}+\cdots+b_{n}y_{n}\in\mathfrak{m}^{2}.$
It follows from the choice of $x_{1},\ldots,x_{j}$ that $a_{1}x_{1}+\cdots+a_{j}x_{j}\in J$, and $b_{1}y_{1}+\cdots+b_{n}y_{n}\in\mathfrak{m}^{2}+J$. Thus, $b_{1},\ldots,b_{n}\in\mathfrak{m}$, since the images of $y_{1},\ldots,y_{n}$ in $\mathfrak{m}/(\mathfrak{m}^{2}+J)$ are linearly independent over $R/\mathfrak{m}$. Then, $b_{1}y_{1}+\cdots b_{n}y_{n}\in\mathfrak{m}^{2}$, and $a_{1}x_{1}+\cdots+a_{j}x_{j}\in\mathfrak{m}^{2}$. Hence, $a_{1},\ldots,a_{j}\in\mathfrak{m}$, since the images of $x_{1},\ldots,x_{j}$ in $\mathfrak{m}/\mathfrak{m}^{2}$ are linearly independent over $R/\mathfrak{m}$. ∎
Theorem 5.2.
Let $(R,\mathfrak{m})$ be an LC-finite regular local ring and $I\supset J$ ideals of $R$ with $\operatorname{depth}_{J}(R)\geq 2$. Suppose that each of $I$ and $J$ is generated by a part of a regular system of parameters. Then, for any $i\geq 1$,
$\operatorname{Ass}H_{I}^{i}(J)\text{ is finite}\Leftrightarrow\operatorname{Ass}H_{I}^{i-1}(R/J)\text{ is finite.}$
Proof.
It is true for $i\leq 2$, see Section 1. Fix $i\geq 3$. As we have seen in Section 3, we may assume $\operatorname{depth}_{I}(R)=i-1\geq 2$. Since $R/J$ is regular, for the cases $\operatorname{depth}_{I}(R/J)=0,1$, the claim holds by Theorem 3.3. Fix $e=\operatorname{depth}_{I}(R/J)\geq 2$. By Lemma 5.1, there are elements $x_{1},\ldots,x_{e-1}\in I$ such that $R/(x_{1},\ldots,x_{e-1},J)$ is regular. The claim also holds for the case $e\geq 2$, by Theorem 4.3. ∎
We apply the theorem to deduce the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$, based on the finiteness of $\operatorname{Ass}H_{I}^{i-1}(R/J)$, which is guaranteed by Marley’s result[Mar]:
Proposition 5.3.
Let $(R,\mathfrak{m})$ be an LC-finite regular local ring of $\dim R\leq 6$, and $I\supset J$ ideals of $R$. Suppose that each of $I$ and $J$ is generated by a part of a regular system of parameters. Then, $\operatorname{Ass}H_{I}^{i}(J)$ is finite for any $i\geq 0$.
Proof.
$H_{I}^{0}(J)$ is finite over $R$, and it has only a finite number of associated primes. If $\operatorname{depth}_{J}(R)=1$, for any $i\geq 0$, $\operatorname{Ass}H_{I}^{i}(J)$ is finite by the LC-finiteness of $R$, which is isomorphic to $J$ as an $R$-module. We may assume $\operatorname{depth}_{J}(R)\geq 2$, and $i>0$. Since $R$ is Cohen-Macaulay, $\operatorname{ht}J=\operatorname{depth}_{J}(R)\geq 2$, and $\dim R/J=\dim R-\operatorname{ht}J\leq 4$. If $\dim R/J\leq 3$, $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite for all $i>0$ by [Mar, Corollary 2.6]. If $\dim R/J=4$, since $\operatorname{ht}I\geq\operatorname{depth}_{J}(R)\geq 2$, $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite for all $i>0$ by [Mar, Proposition 2.7]. Hence, $\operatorname{Ass}H_{I}^{i}(J)$ is finite for any $i>0$, by Theorem 5.2. ∎
Let $(R,\mathfrak{m})$ be a regular local ring. If $\operatorname{char}R=\operatorname{char}R/\mathfrak{m}$, $R$ is said to be of equal characteristic. If $R$ is not of equal characteristic, either
$(\operatorname{char}R,\operatorname{char}R/\mathfrak{m})=(0,p),$
or
$(\operatorname{char}R,\operatorname{char}R/\mathfrak{m})=(p^{n},p)\quad\text{for some $n>1$}.$
In this case, $R$ is said to be of unequal characteristic. Let $(R,\mathfrak{m})$ be a regular local ring of unequal characteristic, and $\operatorname{char}R/\mathfrak{m}=p$. $R$ is called unramified if $p\notin\mathfrak{m}^{2}$. A regular local ring of equal characteristic is conventionally referred to as unramified.
Proposition 5.4.
Let $(R,\mathfrak{m})$ be an LC-finite regular local ring of $\dim R\leq 7$, and $\operatorname{char}R/\mathfrak{m}=p$. Let $I\supset J$ be ideals of $R$. Suppose that each of $I$ and $J$ is generated by a part of a regular system of parameters. Then, $\operatorname{Ass}H_{I}^{i}(J)$ is finite for any $i\geq 0$, if either $p\in J$ or $p\notin\mathfrak{m}^{2}+J$.
Proof.
We may assume $i>0$, $\operatorname{ht}J\geq 2$, and $\dim R/J=\dim R-\operatorname{ht}J\leq 5$. If $\dim R/J\leq 4$, we can prove the claim using the same method as in Proposition 5.3. If $p\in J$ holds, $R/J$ is of equal characteristic; If $p\notin\mathfrak{m}^{2}+J$ holds, $R/J$ is unramified; in conclusion, $R/J$ is unramified. If $\dim R/J=5$, $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite for all $i>0$, by [Mar, Theorem 2.10]. For $i>0$, the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ follows by Theorem 5.2. ∎
Remark 5.5.
We should emphasize in these propositions that it is not required for a regular local ring $R$ to be of equal characteristic. Indeed, if $R$ is of equal characteristic, the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ follows by the Lyubeznik’s results (see [Lyu1] and [Lyu2]), without the assumptions on $R$ and $I$ in the above propositions :
If $R$ is of equal characteristic $p>0$ and $J$ is generated by a part of a regular system of parameters, $\operatorname{Ass}H_{I}^{i}(J)$ is finite for any ideal $I$ of $R$ and any $i\geq 0$ [Lew, Theorem 5.2].
If $R$ is of equal characteristic 0, we show that $\operatorname{Ass}H_{I}^{i}(J)$ is finite for any pair of ideals $I$, $J$, and for any $i\geq 0$. Let $\hat{R}$ be the $\mathfrak{m}$-adic completion. $\hat{R}$ is a Noetherian local ring, and flat over $R$. Then, $H_{I}^{i}(J\hat{R})\simeq H_{I}^{i}(J)\otimes_{R}\hat{R}$. Furthermore, by [Mat, Theorem 23.2(ii)],
[…]
Acknowledgments
The author would like to express their gratitude to Professor Yasunari Nagai for clear guidance on how to proceed with this research and for many helpful discussions.